OpenStudy (anonymous):

Sally can paint a room in 6 hours while Steve can paint the same room in 7 hours. How long will it take if they paint the room together?

OpenStudy (anonymous):

t=t1t2/(t1+t2)

OpenStudy (paxpolaris):

Let x=the hours it would take to paint together

OpenStudy (anonymous):

so how would you set up the equation?? I'm so bad at that part sorry

OpenStudy (saifoo.khan):

Solve for x, get the time for team work.

OpenStudy (paxpolaris):

So, Sally paints how much of the room in x hours?

OpenStudy (saifoo.khan):

\[\frac16+\frac17=\frac 1x\]Sorry did a typo hehe.

OpenStudy (paxpolaris):

x/6... AND Steve would paint x/7 So you can say \[\large \frac x6+\frac x7=1\]

OpenStudy (anonymous):

I got it, thank you!